OpenStudy (anonymous):

If the first step in the solution of the equation -8 - 7x = -5x - 10 is "add 10," then what should the next step be? a.add 5x b.add 7x c.add 8

OpenStudy (lilg132):

the way i do is: -8 -7x = -5x -10 i bring all the variables to one side and the numbers to the other so -8 + 10 = -5x + 7x 2 = 2x 2/2 = 2x / 2 1 = x

OpenStudy (lilg132):

i think the answer is b